# Topologically protected Bell-cat states in a simple spin model

### v2

@@ I Introduction
We analytically confirm that the pair of bound states are well localized in Fock space and closely resemble coherent spin states. Numerical results further demonstrate that these states are separated from the bulk of the spectrum by a relatively large energy gap. As in condensed matter systems, the bound states are protected by an underlying symmetry of the Hamiltonian, making them robust against symmetry-preserving perturbations. These three properties make them ideal target states for the driving protocol. While the resulting BC state remains susceptible to decoherence—particularly from random noise in the driven magnetic field—we find that the state forms rapidly upon entering the topologically nontrivial phase. This suggests that, in practice, the BC state can be realized within a timescale short enough to outpace decoherence.
@@ II Model
$\hat{H}=v\hat{\sigma}_{x}+\frac{2w}{N}\left(\hat{S}_{+}\hat{\sigma}_{-}+\hat{S}_{-}\hat{\sigma}_{+}\right)$ (1)
@@
The usefulness of this basis comes from the fact that the CS model bears a close resemblance to the Su-Schrieffer-Heeger (SSH) model [7] which describes the electronic properties of a 1D polymer with alternating single and double bonds. Under this comparison the Fock state label $n$ is analogous to the unit cell label and $m$ is analogous to the intra-unit cell label. In addition, like the SSH model, the CS model undergoes a topological phase transition from a trivial phase when $v>w$ to a nontrivial phase when $v<w$ which is marked by the presence of a pair of topologically protected bound states with energy $E=0$. However, unlike the SSH model whose bound states only appear at the edges of the polymer, the CS model’s bound states can be centered at any value of $n$. Their protection relies on the chiral symmetry which the Hamiltonian in Eq. (1) obeys and is defined in terms of the chiral operator $\hat{\sigma}_{z}$ and the relation $\hat{\sigma}_{z}\hat{H}\hat{\sigma}_{z}=-\hat{H}$. The chiral symmetry means that $\hat{\sigma}_{z}$ transforms states with energy $E$ to states with energy $-E$ and since the bound states have $E=0$, they are eigenstates of $\hat{\sigma}_{z}$. This also means that each bound state exists entirely in either the CS $\uparrow$ or $\downarrow$ subspace and that any perturbations respecting the chiral symmetry will not destroy them.
In condensed matter systems, topologically protected bound states are well localized in the material. For now we will assume that the same is true for the protected bound states in the Fock space of the CS model and provide evidence for it later. The assumption of localization allows us to make a mean field approximation for the identical spins’ operators by taking their coherent spin expectation values
@@
This approximation labels the states of the identical spins as points on the surface of a spin-$N/2$ Bloch sphere where $\theta$ and $\phi$ are the polar and azimuthal angles, respectively. The mean field approximation of Eq. (1) is
@@
$d_{z}(\theta,\phi)$ $=$ $0$ (4)
Equation (3) resembles a Bloch Hamiltonian of a two-band model describing a periodic lattice under the tight-binding approximation, however, there is one major difference. In a typical Bloch Hamiltonian the vector $\bm{d}$ is parameterized by the quasi-momentum which is $2\pi$-periodic, so it maps points from a circle (1D) or a torus (2D) to, in general, a 3D vector, whereas $\bm{d}$ in Eq. (3) maps points from the mean-field Bloch sphere to a 2D vector ($d_{z}=0$). The azimuthal angle $\phi$ is indeed $2\pi$-periodic, so it plays the role of a quasi-momentum in the CS model, but the polar angle $\theta$ is not $2\pi$-periodic, so it cannot be considered as another quasi-momentum. Instead, it gives the ‘position’ in Fock space from the relation $n=\frac{N}{2}\cos\theta$.
@@
[…]
@@
Figure 1: Geometric representation of the mean-field Hamiltonian. The vector pointing from the origin of the sphere to its surface represents the mean field state of the identical spins while the direction of the vector $\bm{d}$ in the $d_{x}$, $d_{y}$ plane gives the eigenstate of the mean-field Hamiltonian. For a fixed value of $\theta$, as $\phi$ changes by $2\pi$, the sphere vector traces out a circuit (red). The vector $\bm{d}$ follows the same path, but projected directly down onto the $d_{x}$, $d_{y}$ plane. When the projected circuit encompasses the origin of the plane, the system is in the topologically nontrivial phase with winding number $W=1$. The red circle at $\theta=\theta_{1}$ is the boundary circuit between the topologically trivial and nontrivial phases where any circle on the sphere with a larger radius will encompass the origin when projected down onto the plane.
Figure 1 gives a geometric interpretation of the topological properties of the system. The vector extending from the sphere’s center to its surface is $\frac{2}{N}\langle\bm{\hat{S}}\rangle$ (scaled by $w$) and represents the mean field state of the identical spins while the vector in the $d_{x}$, $d_{y}$ plane is $\bm{d}(\theta,\phi)$ whose direction denotes the eigenstate of the mean-field Hamiltonian in Eq. (3). For a fixed value of $\theta$, as $\phi$ changes by $2\pi$, the vector on the sphere traces out a circle on its surface and the vector $\bm{d}(\theta,\phi)$ traces the same path projected directly down onto the $d_{x}$, $d_{y}$ plane. The $d_{x}$ coordinate of the center of the sphere is $v$, whereas the sphere’s radius is $w$. When $v>w$, the projected path fails to enclose the origin of the plane, causing $\bm{d}(\theta,\phi)$ to oscillate without fully winding as $\phi$ changes by $2\pi$. When $v<w$, there exist certain values of $\theta$ for which the trajectory projected onto the plane encircles the origin, causing $\bm{d}(\theta,\phi)$ to wrap around the origin once. The range of $\theta$ exhibiting nonzero winding is $\theta_{1}<\theta<\theta_{2}$, with the boundaries determined by the points where the projected trajectory intersects the origin of the plane. In Figure 1, the red circles represent the trajectories for $\theta=\theta_{1}$ and show that when $\phi=\pi$, $|\bm{d}(\theta_{1},\pi)|=0$ leading to the intersection of the trajectory with the origin and the closure of the energy gap, $E_{\mathrm{gap}}=2|\bm{d}|$, at that point. Although not shown, due to the spherical symmetry, the same thing happens at $(\theta,\phi)=(\theta_{2},\pi)$. Consequently, these points correspond to the locations of the protected bound states.
@@ III.1 Bound state fluctuations
Before describing our protocol for the the creation of BC states we will confirm the existence of the protected bound states discussed in the previous section. The mean field analysis determined the locations of the bound states, however, we would also like to know about their fluctuations to determine how localized they are. For that, we use the fact that the bound states are zero energy eigenstates and exist entirely in one of the CS subspaces, so $\mathcal{H}(n,\phi)\chi_{\uparrow(\downarrow)}(n)=0$ where
@@
The bound states are written as products of the wave function in the identical spins’ Fock space $B_{\uparrow(\downarrow)}(n)$ and a two component vector for the two states of the spin-1/2 particle. Note that we have changed from the polar coordinate, $\theta$, to the $\hat{S}_{z}$ state label, $n$, through the relation $n=\frac{N}{2}\cos\theta$ which will be useful in the coming calculations. In the $n$ state space the bound states are located at $n_{1}=n_{b}$ and $n_{2}=-n_{b}$ where $n_{b}=\frac{N}{2}\sqrt{1-\left(\frac{v}{w}\right)^{2}}$. We are interested in fluctuations around the gap closing point at $\phi=\pi$, so we expand $\mathcal{H}(n,\phi)$ to linear order around this point. In addition, we incorporate fluctuations into the mean-field approximation by making use of the fact that $n$ and $\phi$ are conjugate variables so the transformation $\phi\to i\frac{\partial}{\partial n}$ can be made resulting in the differential equation
@@
$\left.-w\sqrt{1-\left(\frac{2n}{N}\right)^{2}}\left(i\frac{\partial}{\partial n}-\pi\right)\hat{\sigma}_{y}\right]\chi_{\uparrow(\downarrow)}(n)=0\ .$
The wave functions $B_{\uparrow(\downarrow)}(n)$ can be solved for exactly and are of the form $B_{\uparrow}(n)\propto e^{i\pi n}e^{n-\frac{Nv}{2w}\arcsin\left(\frac{2n}{N}\right)}$ where $B_{\downarrow}(n)=B_{\uparrow}(-n)$, however, they are only valid in the region around ±$\pm n_{b}$, so we expand them around those points giving
@@
where $\sigma=\sqrt{\frac{Nv^{2}}{4w^{2}\sqrt{1-\left(\frac{v}{w}\right)^{2}}}}$ is the standard deviation of the wave function. The bound states take the form of Gaussians centered at the gap closing locations with widths that scale as $1/\sqrt{N}$ relative to $N$, so for a sufficiently large number of identical spins, they are well localized. The presence of the $\left[1-(v/w)^{2}\right]^{-1/2}$ factor in $\sigma$ also indicates that the bound states become increasingly localized as the system moves deeper into the topologically nontrivial phase. Indeed, in the limit $v/w\to 0$ the bound states are Fock states located at ±$\pm N/2$. As shown in Figure 1, in this limit, the Bloch sphere sits directly above the origin of the plane, resulting in the bound states being positioned at the North and South poles of the Bloch sphere.
@@ III.2 Generating BC states
Our motivation for starting the process in the trivial phase comes from the fact that eigenstates of the Hamiltonian which will become the topologically protected bound states are centered at $\theta=\pi/2$ ($n=0$ in Fock space) and $\phi=\pi$ on the Bloch sphere in the trivial phase. Furthermore, we start with an equal superposition of the two CS states because we recall that in the nontrivial phase each of the two bound states is in opposite subspaces of the CS. (2) Next, we adiabatically drive the system into the topologically nontrivial phase by linearly ramping down an applied magnetic field which controls the CS spin-flip energy, $v(t)=v_{0}-\gamma t$, where $v_{0}$ is the initial value of $v$ and $\gamma$ is the drive rate. Because the initial state is in a superposition of the CS being both up and down, the wave function will split in two with each part being attached to $B_{\uparrow}(n)$ and $B_{\downarrow}(n)$. (3) After some time when the two parts of the wave function have negligible overlap, they are in a macroscopic superposition of the majority of the identical spins pointing up and pointing down. The splitting process relies on the fact that the topological bound states are protected by the chiral symmetry and are therefore eigenstates of $\hat{\sigma}_{z}$. This forces the bound states on opposite sides of $n=0$ to have support in opposite CS subspaces resulting in the formation of a BC state.
@@
Two primary factors influencing the quality of the final BC state are adiabaticity and the overlap between the initial coherent state and the eigenstates which will become the bound states. In Figure 4 (a), the energy spectrum as a function of $v$ is shown, with the red and black curves representing the energies of the bound and bulk states, respectively. Adiabaticity can be maintained as long as the driving rate is considerably smaller than the minimum gap. While the gap does close between the two $E=0$ eigenstates, they are even and odd superpositions of the two bound states, so they possess opposite parity. As the initial coherent state has even parity under the transformation $n\rightarrow-n$ and flipping of the CS, it has significant overlap with only the even superposition over the course of the dynamics. With adiabaticity satisfied, the fidelity of the final state is determined by the fidelity of the initial state. In the topologically trivial phase, the identical spins’ component of the eventual bound states takes the form of Eq. (8) with $n_{b}=0$ and $\sigma=\sqrt{\frac{N}{4\sqrt{1-\frac{w}{v}}}}$ (see Appendix A for calculation), so when $v\gg w$, the standard deviation is $\sigma\approx\sqrt{N/4}$ and they are approximately equal to the initial coherent state in Eq. (10). Thus, the quality of the final BC state improves as the system is initialized deeper within the topologically trivial phase, which corresponds to larger initial values of $v$. This trend is illustrated in Figure 4 (b), where the final fidelity is displayed to asymptotically approach unity as $v_{0}\to\infty$.
@@
Figure 4: Energy spectrum and final state fidelity. (a) Energy spectrum as a function of the CS spin-flip energy $v$. The red and black curves represent the energies of the bound and bulk states, respectively. The parameters are $N=10$ and $w=1$. (b) Fidelity between the final state at the end of the driving period and the instantaneous energy eigenstate which is a BC state. The parameters are $N=150$, $w=1$, $\gamma\approx 1.2\times 10^{-3}$ and $v_{f}=0.7$.

### v3

@@ I Introduction
We analytically confirm that the pair of bound states are well localized in Fock space and closely resemble coherent spin states. Numerical results further demonstrate that these states are separated from the bulk of the spectrum by a relatively large energy gap. As in condensed matter systems, the bound states are protected by an underlying symmetry of the Hamiltonian, making them robust against symmetry-preserving perturbations. These three properties make them ideal target states for the driving protocol. While the resulting BC state remains susceptible to decoherence, particularly from random noise in the driven magnetic field, we find that the state forms rapidly upon entering the topologically nontrivial phase. This suggests that, in practice, the BC state can be realized within a timescale short enough to outpace decoherence.
@@ II Model
[…]
@@
The usefulness of this basis comes from the fact that the CS model bears a close resemblance to the Su-Schrieffer-Heeger (SSH) model [7] which describes the electronic properties of a 1D polymer with alternating single and double bonds. Under this comparison the Fock state label $n$ is analogous to the unit cell label and $m$ is analogous to the intra-unit cell label. In addition, like the SSH model, the CS model undergoes a topological phase transition from a trivial phase when $v>w$ to a nontrivial phase when $v<w$ which is marked by the presence of a pair of topologically protected bound states with energy $E=0$. However, unlike the SSH model whose bound states only appear at the edges of the polymer, the CS model’s bound states can be centered at any value of $n$. Their protection relies on the chiral symmetry which the Hamiltonian in Eq. (1) obeys and is defined in terms of the chiral operator $\hat{\sigma}_{z}$ and the relation $\hat{\sigma}_{z}\hat{H}\hat{\sigma}_{z}=-\hat{H}$. The chiral symmetry means that $\hat{\sigma}_{z}$ transforms states with energy $E$ to states with energy E-E and since the bound states have $E=0$, they are eigenstates of $\hat{\sigma}_{z}$. This also means that each bound state exists entirely in either the CS $\uparrow$ or $\downarrow$ subspace and that any perturbations respecting the chiral symmetry will not destroy them.
In condensed matter systems, topologically protected bound states are well localized in the material. For now we will assume that the same is true for the protected bound states in the Fock space of the CS model and provide evidence for it later. The assumption of localization allows us to make a mean-field approximation for the identical spins’ operators by taking their coherent spin expectation values
@@
This approximation labels the states of the identical spins as points on the surface of a spin-$N/2$ Bloch sphere where $\theta$ and $\phi$ are the polar and azimuthal angles, respectively. The mean-field approximation of Eq. (1) is
@@
$d_{z}(\theta,\phi)$ $=$ $0.$ (4)
Equation (3) resembles a Bloch Hamiltonian of a two-band model describing a periodic lattice under the tight-binding approximation, however, there is a major difference. In a typical Bloch Hamiltonian the vector $\bm{d}$ is parameterized by the quasi-momentum which is $2\pi$-periodic, so it maps points from a circle (1D) or a torus (2D) to, in general, a 3D vector, whereas $\bm{d}$ in Eq. (3) maps points from the mean-field Bloch sphere to a 2D vector ($d_{z}=0$). The azimuthal angle $\phi$ is indeed $2\pi$-periodic, so it plays the role of a quasi-momentum in the CS model, but the polar angle $\theta$ is not $2\pi$-periodic, so it cannot be considered as another quasi-momentum. Instead, it gives the ‘position’ in Fock space from the relation $n=\frac{N}{2}\cos\theta$.
@@
$W(\theta)$ $=$ $\frac{1}{2\pi}\int_{0}^{2\pi}\left(\bm{\hat{d}}(\theta,\phi)\times\frac{d}{d\phi}\bm{\hat{d}}(\theta,\phi)\right)_{z}d\phi$ (6)
@@
Figure 1: Geometric representation of the mean-field Hamiltonian. The vector pointing from the origin of the sphere to its surface represents the mean-field state of the identical spins while the direction of the vector $\bm{d}$ in the $d_{x}$, $d_{y}$ plane gives the eigenstate of the mean-field Hamiltonian. For a fixed value of $\theta$, as $\phi$ changes by $2\pi$, the sphere vector traces out a circuit (red). The vector $\bm{d}$ follows the same path, but projected directly down onto the $d_{x}$, $d_{y}$ plane. When the projected circuit encompasses the origin of the plane, the system is in the topologically nontrivial phase with winding number $W=1$. The red circle at $\theta=\theta_{1}$ is the boundary circuit between the topologically trivial and nontrivial phases where any circle on the sphere with a larger radius will encompass the origin when projected down onto the plane.
Figure 1 gives a geometric interpretation of the topological properties of the system. The vector extending from the sphere’s center to its surface is $\frac{2}{N}\langle\bm{\hat{S}}\rangle$ (scaled by $w$) and represents the mean-field state of the identical spins while the vector in the $d_{x}$, $d_{y}$ plane is $\bm{d}(\theta,\phi)$ whose direction denotes the eigenstate of the mean-field Hamiltonian in Eq. (3). For a fixed value of $\theta$, as $\phi$ changes by $2\pi$, the vector on the sphere traces out a circle on its surface and the vector $\bm{d}(\theta,\phi)$ traces the same path projected directly down onto the $d_{x}$, $d_{y}$ plane. The $d_{x}$ coordinate of the center of the sphere is $v$, whereas the sphere’s radius is $w$. When $v>w$, the projected path fails to enclose the origin of the plane, causing $\bm{d}(\theta,\phi)$ to oscillate without fully winding as $\phi$ changes by $2\pi$. When $v<w$, there exist certain values of $\theta$ for which the trajectory projected onto the plane encircles the origin, causing $\bm{d}(\theta,\phi)$ to wrap around the origin once. The range of $\theta$ exhibiting nonzero winding is $\theta_{1}<\theta<\theta_{2}$, with the boundaries determined by the points where the projected trajectory intersects the origin of the plane. In Figure 1, the red circles represent the trajectories for $\theta=\theta_{1}$ and show that when $\phi=\pi$, $|\bm{d}(\theta_{1},\pi)|=0$ leading to the intersection of the trajectory with the origin and the closure of the energy gap, $E_{\mathrm{gap}}=2|\bm{d}|$, at that point. Although not shown, due to the spherical symmetry, the same thing happens at $(\theta,\phi)=(\theta_{2},\pi)$. Consequently, these points correspond to the locations of the protected bound states.
@@ III.1 Bound state fluctuations
Before describing our protocol for the the creation of BC states we will confirm the existence of the protected bound states discussed in the previous section. The mean-field analysis determined the locations of the bound states, however, we would also like to know about their fluctuations to determine how localized they are. For that, we use the fact that the bound states are zero energy eigenstates and exist entirely in one of the CS subspaces, so $\mathcal{H}(n,\phi)\chi_{\uparrow(\downarrow)}(n)=0$ where
@@
The bound states are written as products of the wave function in the identical spins’ Fock space $B_{\uparrow(\downarrow)}(n)$ and a two component vector for the two states of the spin-1/2 particle. Note that we have changed from the polar coordinate, $\theta$, to the $\hat{S}_{z}$ state label, $n$, through the relation $n=\frac{N}{2}\cos\theta$ which will be useful in the coming calculations. In the $n$ state space the bound states are located at $n_{1}=n_{b}$ and $n_{2}=-n_{b}$ where $n_{b}=\frac{N}{2}\sqrt{1-\left(\frac{v}{w}\right)^{2}}$. We are interested in fluctuations around the gap closing point at $\phi=\pi$, so we expand $\mathcal{H}(n,\phi)$ to linear order around this point. In addition, we incorporate fluctuations into the mean-field approximation by making use of the fact that the number variable $n$ and the phase variable $\phi$ form a conjugate pair [42, 43] so the transformation $\phi\to i\frac{\partial}{\partial n}$ can be made resulting in the differential equation
@@
$\left.-w\sqrt{1-\left(\frac{2n}{N}\right)^{2}}\left(i\frac{\partial}{\partial n}-\pi\right)\hat{\sigma}_{y}\right]\chi_{\uparrow(\downarrow)}(n)=0\ .$ (8)
The wave functions $B_{\uparrow(\downarrow)}(n)$ can be solved for exactly and are of the form $B_{\uparrow}(n)\propto e^{i\pi n}e^{n-\frac{Nv}{2w}\arcsin\left(\frac{2n}{N}\right)}$ where $B_{\downarrow}(n)=B_{\uparrow}(-n)$, however, they are only valid in the region around $\pm n_{b}$, so we expand them around those points giving
@@
where $\sigma=\sqrt{\frac{Nv^{2}}{4w^{2}\sqrt{1-\left(\frac{v}{w}\right)^{2}}}}$ is the standard deviation of the wave function. The bound states take the form of Gaussians centered at the gap closing locations with widths that scale as $1/\sqrt{N}$ relative to $N$, so for a sufficiently large number of identical spins, they are well localized. The presence of the $\left[1-(v/w)^{2}\right]^{-1/2}$ factor in $\sigma$ also indicates that the bound states become increasingly localized as the system moves deeper into the topologically nontrivial phase. Indeed, in the limit $v/w\to 0$ the bound states are Fock states located at $\pm N/2$. As shown in Figure 1, in this limit, the Bloch sphere sits directly above the origin of the plane, resulting in the bound states being positioned at the north and south poles of the Bloch sphere.
@@ III.2 Generating BC states
Our motivation for starting the process in the trivial phase comes from the fact that eigenstates of the Hamiltonian which will become the topologically protected bound states are centered at $\theta=\pi/2$ ($n=0$ in Fock space) and $\phi=\pi$ on the Bloch sphere in the trivial phase. Furthermore, we start with an equal superposition of the two CS states because we recall that in the nontrivial phase each of the two bound states is in opposite subspaces of the CS. (2) Next, we adiabatically drive the system into the topologically nontrivial phase by linearly ramping down an applied magnetic field which controls the CS spin-flip energy,
$v(t)=v_{0}-\gamma t,$ (12)
where $v_{0}$ is the initial value of $v$ and $\gamma$ is the drive rate. Because the initial state is in a superposition of the CS being both up and down, the wave function will split in two with each part being attached to $B_{\uparrow}(n)$ and $B_{\downarrow}(n)$. (3) After some time when the two parts of the wave function have negligible overlap, they are in a macroscopic superposition of the majority of the identical spins pointing up and pointing down. The splitting process relies on the fact that the topological bound states are protected by the chiral symmetry and are therefore eigenstates of $\hat{\sigma}_{z}$. This forces the bound states on opposite sides of $n=0$ to have support in opposite CS subspaces resulting in the formation of a BC state.
@@
Two primary factors influencing the quality of the final BC state are adiabaticity and the overlap between the initial coherent state and the energy eigenstate which will become the bound states. To ensure adiabaticity, we require that the drive rate is small enough to satisfy [44],
$\max\limits_{v_{0}\leq v\leq v_{f}}\frac{|\langle E_{n}(v)|\dot{\hat{H}}|E_{m}(v)\rangle|}{|E_{n}(v)-E_{m}(v)|^{2}}\ll 1,\hskip 15.0pt\text{for all}\hskip 5.0ptm\neq n$ (14)
where $v_{f}$ is the final value of $v$ and $|E_{n}(v)\rangle$ is the $n^{\mathrm{th}}$ instantaneous energy eigenstate for a given value of $v$. Equation 14 can be rearranged in terms of the drive rate which results in the condition $\gamma\ll\gamma^{\prime}$ where, for the CS model,
$\gamma^{\prime}=\min\limits_{v_{0}\leq v\leq v_{f}}\frac{|E_{N+2}(v)-E_{m}(v)|^{2}}{|\langle E_{N+2}(v)|\hat{\sigma}_{x}|E_{m}(v)\rangle|},\hskip 5.0pt\text{for}\hskip 5.0ptm\neq N+2,$ (15)
where $|E_{N+2}(v)\rangle$ is the target BC state in the topological phase and we have used the fact that $\dot{\hat{H}}=-\gamma\hat{\sigma}_{x}$. In Figure 4 (a), the energy spectrum as a function of $v$ is shown, with the red and black curves representing the energies of the bound and bulk states, respectively. The energy of the target state is the top red curve in the image. Although the gap between the two $E=0$ eigenstates closes, each eigenstate is an opposite parity superposition of the two bound states. Because the initial state in Eq. (10) has even parity and the driven operator $\hat{\sigma}_{x}$ does not couple states of opposite parity, $\gamma^{\prime}$ stays nonzero throughout the driving period for any finite system size. We find that $\gamma^{\prime}$ is a minimum near the critical point ($v=w$) and corresponds to the adjacent even parity eigenstate ($m=N+4$). Numerically, we find $\gamma^{\prime}\approx 22.35N^{-1.33}$ (see Appendix A) suggesting a finite size scaling exponent of 4/3. For the system sizes used in this work, $100\leq N\leq 200$, this corresponds to $1.9\times 10^{-2}\leq\gamma^{\prime}\leq 4.8\times 10^{-2}$. Throughout the paper we use a drive rate of $\gamma=1.2\times 10^{-3}$ which satisfies the adiabaticity condition.
@@
With adiabaticity satisfied, the fidelity of the final state is determined by the fidelity of the initial state. In the topologically trivial phase, the identical spins’ component of the eventual bound states takes the form of Eq. (9) with $n_{b}=0$ and $\sigma=\sqrt{\frac{N}{4\sqrt{1-\frac{w}{v}}}}$ (see Appendix B for calculation), so when $v\gg w$, the standard deviation is $\sigma\approx\sqrt{N/4}$ and they are approximately equal to the initial coherent state in Eq. (11). Thus, the quality of the final BC state improves as the system is initialized deeper within the topologically trivial phase. This trend is illustrated in Figure 4 (b), where the final fidelity is displayed to asymptotically approach unity as $v_{0}\to\infty$. Numerical results of the finite size scaling of the initial fidelity can be found in Appendix A.
We now briefly discuss the effects of chiral symmetry breaking terms such as $v_{z}\hat{\sigma}_{z}$ and $2w_{z}\hat{\sigma}_{z}\hat{S}_{z}/N$, and estimate the allowable strengths of these perturbations for the BC state to remain generable. Both terms shift the energies of the even and odd parity superpositions of the bound states away from $E=0$, but in different ways. The first term breaks parity symmetry by destroying the invariance of the Hamiltonian under the combined transformation $\hat{\sigma}_{z}\to-\hat{\sigma}_{z}$ and $\hat{S}_{z}\to-\hat{S}_{z}$, leading to an energy splitting of $\pm E_{v_{z}}=\pm v_{z}$ to leading order for small $v_{z}$. The second term preserves parity symmetry, but the factor of $\hat{S}_{z}$ introduces a higher energy cost for eigenstates with components farther away from $n=0$, i.e., BC states whose superposition has larger separation. This term increases the energies of both parity eigenstates by the same amount, and for small $w_{z}$ we estimate the shift to be $E_{w_{z}}=2w_{z}n_{b}/N$, where $n_{b}=N\sqrt{1-(v/w)^{2}}/2$ is the mean-field magnitude of the bound state location calculated in the previous subsection.
The chiral symmetry breaking perturbations will reduce the energy gap, $E_{\mathrm{gap}}$, between the bound and bulk states by a small amount, which in turn slightly decreases $\gamma^{\prime}$. This reduction can be compensated by slowing the drive rate, and as long as the gap remains open, the BC state can still be generated. Therefore, we require the shifts $E_{v_{z}}$ and $E_{w_{z}}$ to satisfy $E_{v_{z}},E_{w_{z}}\ll E_{\mathrm{gap}}$. To estimate $E_{\mathrm{gap}}$, we first note that in the topological phase ($v/w<1$) the energy spectrum is approximately independent of $v$, as seen in Fig. 4(a). This allows us to set $v=0$, leaving only the spin-exchange term in $\hat{H}$. In this limit, we are able to diagonalize $\hat{H}$ in the subspace spanned by $|n,\uparrow\rangle$ and $|n+1,\downarrow\rangle$ for each $n$ giving energies $E_{n,\pm}=\pm 2w\sqrt{\left(N/2-n\right)\left(N/2+n+1\right)}/N$. The energy gap is then $E_{\mathrm{gap}}\approx E_{N/2-1,+}-E_{N/2,+}=2w/\sqrt{N}$. Thus, the BC state remains generable provided $v_{z}\ll 2w/\sqrt{N}$ and $w_{z}\ll 2w/\sqrt{N\left[1-(v/w)^{2}\right]}$.
@@
Figure 4: Energy spectrum and final state fidelity. (a) Energy spectrum as a function of the scaled CS spin-flip energy $v/w$. The red and black curves represent the energies of the bound and bulk states, respectively. The parameters are $N=10$ and $w=1$. (b) Fidelity between the final state at the end of the driving period and the BC state as a function of the scaled initial CS spin-flip energy $v_{0}/w$. The parameters are $N=150$, $w=1$, $\gamma\approx 1.2\times 10^{-3}$ and $v_{f}=0.7$.